Liddle syndrome is a rare hereditary disorder in which the kidneys excrete K, but retain too much Na and water, leading to hypertension (high blood pressure). This syndrome is autosomal dominant, meaning that the presence of a single allele for the syndrome will result in the disease. Patients usually show symptoms before they are 35 years of age. A woman with Liddle syndrome has a child with a normal male. What is the probability that their child will have the syndrome?
Answer #1

Let the allele which causes disease to be L and recessive allele be l then

Female can be LL or Ll ( since the disease the autosomal dominant)

and the male is normal so he is ll

1) woman is LL

L L
l Ll Ll
l Ll Ll

here all the progenies have L allele so

Probability of getting an affected child= number of affected child/total number

=4/4

=1

2) woman is Ll

L l
l Ll ( affected) ll (normal)
l Ll (affected) ll (normal)

Probability of getting an affected child= number of affected child/total number

=2/4

=1/2